OpenStudy (anonymous):

A cookbook divides the following general guidelines for use of a pressure cooker. "For steaming vegetables,cooking time at a gauge pressure of 15 pounds per square inch (psi) is 1/3 that at atmospheric pressure." Remember that gauge pressure is measure relative to the external atmospheric pressure, which is 15 psi at sea level. From this info. estimate the activation energy for the process of steaming vegetables. (Hint: Clausius and clapeyron may help) Ans: 64kj/mol

OpenStudy (aaronq):

use the Clausius-Clayperion equation \(ln(\dfrac{P_2}{P_1})=\dfrac{-\Delta H_{vap}}{R}(\dfrac{1}{T_2}-\dfrac{1}{T_1})\)